### v1

# Convergence Rates for Stochastic Proximal and Projection Estimators

Abstract.
In this paper, we establish explicit, non-asymptotic convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [5, 6]. In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections. We prove a dimension-explicit $\sqrt{\delta}$ bound in the $\rho$-weakly convex (possibly nonsmooth) setting and show, by examples, that this order is sharp. Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary.
Key words and phrases:
proximal mapping, weakly convex functions, stochastic approximation, convergence rates
2020 Mathematics Subject Classification:
Primary 49J53; Secondary 49J42, 90C30, 90C25
1. Introduction
Proximal mappings and metric projections are fundamental primitives in variational analysis and optimization. Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$. When $f=\iota_{C}$ is the indicator of a closed convex set $C$, the proximal mapping reduces to the metric projection $\operatorname{proj}_{C}(x)$. In many applications, however, evaluating $\operatorname{prox}_{\lambda f}(x)$ or $\operatorname{proj}_{C}(x)$ exactly is computationally expensive, and in black-box settings only function values (possibly noisy) are available.
This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting. For $\delta>0$, define the barycenter
(1) $m_{\delta}(x):=\frac{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[Y\,\exp(-f(Y)/\delta)\big]}{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[\exp(-f(Y)/\delta)\big]}.$
When $f=\iota_{C}$, the estimator becomes the conditional mean
$p_{\delta}(x)=\mathbb{E}\big[Y\mid Y\in C\big],\qquad Y\sim\mathcal{N}(x,\delta I),$
which can be interpreted as a “smoothed projection” onto $C$.
The estimators (1) arise naturally from Laplace-type asymptotics for integrals with kernel $\exp(-g/\delta)$, where $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$. Two recent strands of work developed this idea.
First, motivated by Hamilton-Jacobi equations, [5] proposed the HJ-Prox method and proved that, for $\rho$-weakly convex objectives under mild regularity, the barycenter converges to the proximal point as $\delta\to 0^{+}$. Second, [6] places the relationship between infimal convolutions and self-normalized Laplace approximation front and center, extending the construction beyond the quadratic kernel and establishing asymptotic validity under weak local assumptions; in particular, their framework includes smoothed set projections $p_{\delta}(x)=\mathbb{E}[Y\mid Y\in C]$ and proves $p_{\delta(x)}\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$ under mild local geometric conditions near the projection point.
While these references provide a compelling unifying viewpoint and broad asymptotic consistency results, they leave open a basic quantitative question that is decisive for algorithmic use: how fast do these stochastic estimators converge, and how do the constants depend on the ambient dimension and on the geometry/smoothness of the problem?
Our contributions: explicit and sharp convergence rates
The main goal of this work is to provide non-asymptotic, dimension-explicit convergence rates for (1) and its projection specialization, and to identify regimes where improved rates hold.
(i) Nonsmooth weakly convex case: a sharp $\sqrt{\delta}$ rate. Assume $f$ is proper, lower semicontinuous, and $\rho$-weakly convex, and fix $\lambda\in(0,1/\rho)$ with $\mu:=\frac{1}{\lambda}-\rho>0$. Under the mild geometric condition that $\operatorname{dom}f$ has nonempty interior, we prove the global bound (valid for every $\delta>0$)
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
This estimate is complemented by explicit examples showing that the $\sqrt{\delta}$ order cannot be improved in general.
(ii) Smooth case: an $O(\delta)$ rate with explicit constants. If $f\in C^{2}(\mathbb{R}^{n})$ is $\rho$-weakly convex and $\nabla^{2}f$ is globally Lipschitz with constant $L$, we establish the linear-rate refinement
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
again for every $\delta>0$. This result quantifies a phenomenon suggested by the smoothing interpretations in [5, 6]: additional curvature regularity improves the bias of the barycentric estimator.
(iii) Projection corollaries and boundary-sensitive refinements. Taking $f=\iota_{C}$ and $\lambda=1$ yields a quantitative convergence estimate for convex projections:
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
Beyond the convex Lipschitz regime, we analyze the case of convex sets with $C^{2,1}$ boundary charts at the projection point and obtain a refined $O(\delta)$ expansion. This provides a bridge between the purely asymptotic projection consistency statement in [6] and the boundary-geometry-dependent expansions needed in fine asymptotic analysis.
The remainder of the paper is organized as follows. Section 2 collects the mathematical preliminaries used throughout. Section 3 establishes the $\sqrt{\delta}$ bound for nonsmooth weakly convex functions, provides sharpness examples, and derives the corollary for convex projections. Section 4 proves the $O(\delta)$ rate under $C^{2}$ regularity with Lipschitz continuous Hessian and develops a refined analysis for convex sets with $C^{2,1}$ boundaries. The paper ends with some concluding remarks and possible lines of future research.
2. Mathematical Preliminaries
Throughout this paper, we work in the finite-dimensional Hilbert space $\mathbb{R}^{n}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ and the associated norm $\|\cdot\|$. The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$. We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
For a ser $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
$\iota_{A}(x):=\left\{\begin{array}[]{cc}0&\text{ if }x\in A,\\ +\infty&\text{ if }x\notin A.\end{array}\right$.
The distance function from $x\in\mathbb{R}^{n}$ to $A$ is given by
$d_{A}(x)=\inf\left\{\|x-y\|:y\in A\right\}.$
Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$. Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{\mathbb{R}}\cup\{+\infty\}$, the domain of $g$ is
$\operatorname{dom}(g):=\{x\in\mathbb{R}^{n}:g(x)<+\infty\}.$
We say that $g$ is proper if $\operatorname{dom}g\neq\emptyset$. Given $\rho\geq 0$, a function $g\colon\mathbb{R}^{n}\to\mathbb{\mathbb{R}}\cup\{+\infty\}$ is called $\rho$-weakly convex if $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. For $\mu>0$, we say that $g$ is $\mu$-strongly convex if $g-\frac{\mu}{2}\|\cdot\|^{2}$ is convex.
The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to{\mathbb{R}}_{\infty}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
$\operatorname{e}_{\lambda}g(x):=\inf_{y\in\mathbb{R}^{n}}\left(g(y)+\frac{1}{2\lambda}\|x-y\|^{2}\right)\textrm{ for all }x\in\mathbb{R}^{n}.$
If $g$ is proper, lower semicontinuous, and $\rho$-weakly convex, and if $0<\lambda<\frac{1}{\rho}$ (with the convention $1/0=+\infty$), then the above infimum is attained at a unique point $\operatorname{prox}_{\lambda g}(x)\in\mathbb{R}^{n}$, and
$\operatorname{e}_{\lambda}g(x)$ $=g(\operatorname{prox}_{\lambda g}(x))+\frac{1}{2\lambda}\|x-\operatorname{prox}_{\lambda g}(x)\|^{2}.$
In this case, the operator $x\mapsto\operatorname{prox}_{\lambda g}(x)$ is everywhere defined and is called the proximal operator of $g$ of index $\lambda$. Moreover, $\operatorname{e}_{\lambda}g$ is continuously differentiable and, for each $x\in\mathbb{R}^{n}$,
$\nabla\operatorname{e}_{\lambda}g(x)=\frac{1}{\lambda}(x-\operatorname{prox}_{\lambda g}(x)).$
We refer to [1, 2] for more details.
Let $\delta>0$, and let $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a proper function such that $e^{-g/\delta}$ is integrable and
$\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy>0.$
Under these assumptions, $e^{-g/\delta}$ induces a probability measure. We denote by
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
the expectation of a given function $w$ with respect to this probability measure.
The following result is known as Fradelizi’s inequality (see [4])
Lemma 2.1.
Let $h:\mathbb{R}^{n}\to[0,\infty)$ be an integrable log-concave function, and let
$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}$
denotes its barycenter. Then
$h(b_{h})\ \geq\ e^{-n}\,\sup_{z\in\mathbb{R}^{n}}h(z).$
3. Nonsmooth results
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The aim of this section is to prove that, for any $0<\lambda<1/\rho$,
$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[y\cdot\exp(-f(y)/\delta)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[\exp(-f(y)/\delta)]}\to\operatorname{prox}_{\lambda f}(x)\textrm{ as }\delta\to 0^{+}.$
The next result provides a rate of convergence of the barycenter toward the projection.
Theorem 3.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper, lower semicontinuous, and $\rho$-weakly convex for some $\rho\geq 0$. Assume that $\operatorname{dom}f$ has nonempty interior. Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex, it follows that $g$ is $\mu$-strongly convex and
(2) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
Consider the vector field
$F(y):=(y-\operatorname{prox}_{\lambda f}(x))e^{-g(y)/\delta}.$
Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta$.
Proof of Claim 1: For $R>0$, apply the divergence theorem on $\mathbb{B}_{R}$:
$\int_{\mathbb{B}_{R}}\operatorname{div}F(y)dy=\int_{\partial\mathbb{B}_{R}}\langle F(y),\nu(y)\rangle dS(y).$
Using (2), the right-hand side is bounded in absolutely value by
$CR^{n}\exp\left(-\frac{\mu}{2\delta}(R-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }R\to\infty,$
hence $\int_{\mathbb{R}^{n}}\operatorname{div}F(y)dy=0$. Since $g$ is convex, it is differentiable a.e. and
$\operatorname{div}F(y)dy=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}\textrm{ for a.e. }y\in\mathbb{R}^{n}.$
Finally, integrating over $\mathbb{R}^{n}$ and dividing by $\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy$, we obtain the claim.
Next, by strong convexity and almost-everywhere differentiability, we obtain
$\langle\nabla g(y),y-\operatorname{prox}_{\lambda f}(x)\rangle\geq\mu\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}.$
Taking expectation yields
$\mu\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta,$
hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen inequality,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}$ $\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|\right]^{2}$
$\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]$
$\leq\frac{n\delta}{\mu}.$
Remark 3.2.
Let $Y\sim\sigma_{\delta}$, where $\sigma_{\delta}$ has density proportional to $\exp(-g/\delta)$. The proof of Theorem 3.1 yields the mean-square localization estimate
$\mathbb{E}\big[\|Y-\operatorname{prox}_{\lambda f}(x)\|^{2}\big]\leq\frac{n\delta}{\mu}.$
This bound also yields a simple concentration inequality: for every $r>0$,
$\sigma_{\delta}\!\big(\|Y-\operatorname{prox}_{\lambda f}(x)\|\geq r\big)\leq\frac{n\delta}{\mu\,r^{2}},$
and therefore, for any $\eta\in(0,1)$,
$\sigma_{\delta}\!\left(B_{\sqrt{n\delta/(\mu\eta)}}(\operatorname{prox}_{\lambda f}(x))\right)\geq 1-\eta.$
The following two examples show that our convergence rates are sharp.
Example 3.1.
Fix $n\geq 1$ and consider
$f(y)=\sum_{i=1}^{n}\max\{y_{i},0\}.$
Then $f$ is proper, lower semicontinuous, convex and $\operatorname{dom}f=\mathbb{R}^{n}$ has nonempty interior. Fix $x=0$ and any $\lambda>0$, and set $\mu=1/\lambda$. Then, $\operatorname{prox}_{\lambda f}(0)=0$ and, by symmetry of $f$, for any $\delta>0$
$m_{\delta}(0)=(a_{\delta},\ldots,a_{\delta}),$
where
$a_{\delta}:=\frac{\int_{\mathbb{R}}t\,e^{-\phi(t)/\delta}\,dt}{\int_{\mathbb{R}}e^{-\phi(t)/\delta}\,dt}\qquad\phi(t):=\max\{t,0\}+\frac{1}{2\lambda}t^{2}.$
Moreover, one can show that
$a_{\delta}=-\sqrt{\frac{2}{\pi}\lambda\delta}+O(\delta)\textrm{ as }\delta\to 0^{+}.$
Consequently,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\sqrt{n}|a_{\delta}|=\sqrt{\frac{2n}{\pi\mu}\delta}+O(\delta).$
This shows that the order of convergence in Theorem 3.1 is sharp.
The following example shows that the bound obtained in Theorem 3.1 is sharp.
Example 3.2.
Fix $n\geq 2$, take $\lambda>0$, set $\mu=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\}.$
Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]m_{\alpha,n},$
where $\chi_{n}$ denotes the $\chi$-distribution with $n$ degrees of freedom and
$m_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$. In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $m_{\alpha,n}\in[\cos\alpha,1]$ and hence $m_{\alpha,n}\to 1$ as $\alpha\searrow 0$. Furthermore,
$\mathbb{E}[\chi_{n}]\sqrt{n}\to 1\textrm{ as }n\to\infty.$
Consequently, for every $\varepsilon>0$ there exist $n$ and $\alpha$ such that, for all $\delta>0$,
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
This shows that the rate of convergence in Theorem 3.1 is sharp.
We end this section by showing that $m_{\delta}(x)$ coincides with $\operatorname{prox}_{\lambda f}(x)$ whenever $f$ is quadratic.
Example 3.3.
Let $A$ be a symmetric and satisfy $A\succeq-\rho I$ for some $\rho\geq 0$, and define
$f(y)=\frac{1}{2}\langle Ay,y\rangle+\langle b,y\rangle+c.$
Then for every $x\in\mathbb{R}^{n}$, every $\delta>0$, and every $0<\lambda<1/\rho$ (with the convention $1/0=+\infty$), one has
$m_{\delta}(x)=\operatorname{prox}_{\lambda f}(x)=\left(A+\frac{1}{\lambda}I\right)^{-1}\left(\frac{1}{\lambda}x-b\right).$
3.1. Stochastic Approximation of Projections onto Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. We define
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
As a direct consequence of Theorem 3.1, we obtain a quantified convergence result: $p_{\delta}(x)\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$.
Corollary 3.3.
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Fix $x\in\mathbb{R}^{n}$. Then, for every $\delta>0$,
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
4. Smooth results
In this section, we show that additional smoothness of the problem data yields improved convergence rates compared with those in Section 3.
Theorem 4.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}$ be $C^{2}$ and $\rho$-weakly convex with $\rho\geq 0$. Assume that $\nabla^{2}f$ is globally Lipschitz: there exists $L\geq 0$ such that
$\|\nabla^{2}f(u)-\nabla^{2}f(v)\|\leq L\|u-v\|\quad\textrm{ for all }u,v\in\mathbb{R}^{n}.$
Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Let us consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and
(3) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
From this inequality we get that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
Given a function $w(\cdot)$, we define
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
Claim 1: $\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=0$.
Proof of Claim 1: We observe that
$\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)=-\frac{1}{\delta}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}.$
Moreover, from (3), it follows that
$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
which implies that
$0=\int_{\mathbb{R}^{n}}\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)dy=-\frac{1}{\delta}\int_{\mathbb{R}^{n}}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}dy,$
which implies the claim.
Claim 2: Let $H=\nabla^{2}g(\operatorname{prox}_{\lambda f}(x))$. Then, for all $y\in\mathbb{R}^{n}$
$\nabla g(y)=H(y-\operatorname{prox}_{\lambda f}(x))+r(y),$
where $\|r(y)\|\leq\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}$.
Proof of Claim 2: Since $\nabla g(\operatorname{prox}_{\lambda f}(x))=0$, the formula follows from Taylor’s theorem with
$r(y)=\int_{0}^{1}(\nabla^{2}g(\operatorname{prox}_{\lambda f}(x)+t(y-\operatorname{prox}_{\lambda f}(x)))-H)dt.$
Because $\nabla^{2}g=\nabla^{2}f+\frac{1}{\lambda}I$ and $\nabla^{2}f$ is $L$-Lipschitz, we have
$\|\nabla^{2}g(u)-\nabla^{2}g(v)\|\leq L\|u-v\|.$
Hence
$\|r(y)\|\leq\int_{0}^{1}Lt\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}=\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2},$
which proves the claim.
Claim 3: $m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)=-H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]$.
Proof of Claim 3: From Claim 1 and 2 we have that
$0$ $=\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=H(m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x))+\mathbb{E}_{\sigma_{\delta}}[r(y)],$
which implies the result.
Claim 4: For every $\delta>0$, we have
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\delta.$
Proof of Claim 4: On the one hand, from Claims 2 and 3, we have
(4) $\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|$ $=\|H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]\|$
$\leq\frac{1}{\mu}\mathbb{E}_{\sigma_{\delta}}\left[\|r(y)\|\right]$
$\leq\frac{L}{2\mu}\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}],$
where we have used that $H\succcurlyeq\mu I$ and $\|H^{-1}\|\leq 1/\mu$. Moreover,
(5) $\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]=\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]+\mathbb{E}_{\sigma_{\delta}}[\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}].$
On the other hand, we observe that $g/\delta$ is $\mu/\delta$ strongly convex. Hence, $h_{\delta}=e^{-g/\delta}$ is strongly log-concave. From Brascamp–Lieb inequality (see [3, Theorem 4.1]), we get that
(6) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\quad\Rightarrow\quad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu}.$
Next, again since $h_{\delta}=e^{-g/\delta}$ is log-concave, applying the Lemma 2.1 to $h_{\delta}$ yields
(7) $g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\leq n\delta.$
By $\mu$-strong convexity of $g$,
$g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\geq\frac{\mu}{2}\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2},$
and combining with (7) gives
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{2n\delta}{\mu}.$
Plugging (6) and the above inequality into (5) yields
$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{2n\delta}{\mu}=\frac{3n\delta}{\mu}.$
Finally, from the above inequality and (4), we get
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{3n\delta}{\mu}=\frac{3nL}{2\mu^{2}}\delta,$
which proves Claim 4 and the theorem. ∎
4.1. Stochastic Approximation of Projections onto Smooth Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Recall that
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
Definition 4.2.
Let $C\subset\mathbb{R}^{n}$ be a closed set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
$C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)=\{(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}:\ t\leq h(z)\},$
and $\nabla^{2}h$ is Lipschitz on $\mathbb{B}^{n-1}_{\rho}$ with constant $M\geq 0$, i.e.,
$\|\nabla^{2}h(z)-\nabla^{2}h(z^{\prime})\|\leq M\|z-z^{\prime}\|\qquad\text{for all }z,z^{\prime}\in\mathbb{B}^{n-1}_{\rho}.$
In the case of a smooth convex set, we obtain the following improvement of Corollary 3.3, which provides the convergence rate
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)\textrm{ as }\delta\to 0^{+}.$
Proposition 4.3.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exists $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $K_{n}\delta,$
where $K_{n}$ is a constant depending on the dimension $n$, $\rho$, $K$, and $M$.
Example 4.1.
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0$, where $\|n\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,n\rangle>0$. Then
$\operatorname{proj}_{C}(x)=x-\langle x,n\rangle n,\quad d_{C}(x)=\langle x,n\rangle.$
Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the tangential conditional mean is unchanged and the drift is purely normal. Hence,
$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,n$
for some scalar $\alpha_{\delta}<0$. In fact, one can show that
$\alpha_{\delta}:=d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})},$
and therefore
$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)n,$
which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Proposition 4.3 is sharp.