### v3

# Convergence Rates for Stochastic Proximal and Projection Estimators

Abstract.
In this paper, we establish explicit convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [10, 11]. In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections. We prove a dimension-explicit $\sqrt{\delta}$ bound, with explicit constants for the proximal mapping, in the $\rho$-weakly convex (possibly nonsmooth) setting, and we also obtain a dimension-explicit $\sqrt{\delta}$ rate for the metric projection onto an arbitrary convex set with nonempty interior. Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary. Examples demonstrate that these rates are optimal.
Key words and phrases:
proximal mapping, weakly convex functions, stochastic approximation, convergence rates
2020 Mathematics Subject Classification:
Primary 49J53; Secondary 49J42, 90C30, 90C25
1. Introduction
Proximal mappings and metric projections are fundamental concepts in variational analysis and optimization. Introduced by J.-J. Moreau in [9], proximal mappings have since become central tools in modern optimization, notably through proximal algorithms and operator-splitting methods and their many variants (see, e.g., [2]). Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$. When $f=\iota_{C}$ is the indicator of a closed convex set $C$, the proximal mapping reduces to the metric projection $\operatorname{proj}_{C}(x)$. In many applications, however, evaluating $\operatorname{prox}_{\lambda f}(x)$ or $\operatorname{proj}_{C}(x)$ exactly is computationally expensive, and in black-box settings only function values (possibly noisy) are available.
This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting. For $\delta>0$ and $\lambda>0$, we define the barycenter
(1) $m_{\delta}(x):=\frac{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[Y\,\exp(-f(Y)/\delta)\big]}{\mathbb{E}_{Y\sim\mathcal{N}(x,\delta\lambda I)}\!\big[\exp(-f(Y)/\delta)\big]}.$
When $f=\iota_{C}$, the estimator becomes the conditional mean
$p_{\delta}(x)=\mathbb{E}\big[Y\mid Y\in C\big],\qquad Y\sim\mathcal{N}(x,\delta I),$
which can be interpreted as a “smoothed projection” onto $C$.
The estimators (1) arise naturally from Laplace-type asymptotics for integrals with kernel $\exp(-g/\delta)$, where $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$. Two recent strands of work developed this idea. First, motivated by Hamilton-Jacobi equations, [10] proposed the HJ-Prox method and proved that, for $\rho$-weakly convex objectives under mild regularity, the barycenter converges to the proximal point as $\delta\to 0^{+}$. Second, [11] places the relationship between infimal convolutions and self-normalized Laplace approximation front and center, extending the construction beyond the quadratic kernel and establishing asymptotic validity under weak local assumptions; in particular, their framework includes smoothed set projections $p_{\delta}(x)=\mathbb{E}[Y\mid Y\in C]$ and proves $p_{\delta(x)}\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$ under mild local geometric conditions near the projection point.
While these references provide a compelling unifying viewpoint and broad asymptotic consistency results, they leave open a basic quantitative question that is decisive for algorithmic use: how fast do these stochastic estimators converge, and how do the constants depend on the ambient dimension and on the geometry/smoothness of the problem?
Our contributions: explicit and sharp convergence rates
The main goal of this work is to provide non-asymptotic, dimension-explicit convergence rates for (1) and its projection specialization, and to identify regimes where improved rates hold.
(i) Nonsmooth weakly convex case: a sharp $\sqrt{\delta}$ rate. Assume $f$ is proper, lower semicontinuous, and $\rho$-weakly convex, and fix $\lambda\in(0,1/\rho)$ with $\mu:=\frac{1}{\lambda}-\rho>0$. Under the mild geometric condition that $\operatorname{dom}f$ has nonempty interior, we prove the global bound (valid for every $\delta>0$)
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
This estimate is complemented by explicit examples showing that the $\sqrt{\delta}$ order cannot be improved in general.
(ii) Smooth case: an $O(\delta)$ rate with explicit constants. If $f\in C^{2}(\mathbb{R}^{n})$ is $\rho$-weakly convex and $\nabla^{2}f$ is globally Lipschitz with constant $L$, we establish the linear-rate refinement
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
again for every $\delta>0$. This result quantifies a phenomenon suggested by the smoothing interpretations in [10, 11]: additional curvature regularity improves the bias of the barycentric estimator.
(iii) Projection corollaries and boundary-sensitive refinements. Taking $f=\iota_{C}$ and $\lambda=1$ yields a quantitative convergence estimate for convex projections:
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
Beyond the convex Lipschitz regime, we analyze the case of convex sets with $C^{2,1}$ boundary charts at the projection point and obtain a refined $O(\delta)$ expansion. This provides a bridge between the purely asymptotic projection consistency statement in [11] and the boundary-geometry-dependent expansions needed in fine asymptotic analysis.
The remainder of the paper is organized as follows. Section 2 collects the mathematical preliminaries used throughout. Section 3 establishes the $\sqrt{\delta}$ bound for nonsmooth weakly convex functions, provides sharpness examples, and derives the corollary for convex projections. Section 4 proves the $O(\delta)$ rate under $C^{2}$ regularity with Lipschitz continuous Hessian and develops a refined analysis for convex sets with $C^{2,1}$ boundaries. The paper ends with some concluding remarks and possible lines of future research.
2. Mathematical Preliminaries
Throughout this paper, we work in the finite-dimensional Hilbert space $\mathbb{R}^{n}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ and the associated norm $\|\cdot\|$. The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$. The open ball of radius $R$ centered at $x$ is denoted by $B_{R}(x)$. For simplicity, the open ball of radius $R$ centered at 0 is denoted by $B_{R}$. Moreover, $\mathbb{S}^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
For a set $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
$\iota_{A}(x):=\left\{\begin{array}[]{cc}0&\text{ if }x\in A,\\ +\infty&\text{ if }x\notin A.\end{array}\right$.
The distance function from $x\in\mathbb{R}^{n}$ to $A$ is given by
$d_{A}(x)=\inf\left\{\|x-y\|:y\in A\right\}.$
Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$. Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$, the domain of $g$ is
$\operatorname{dom}(g):=\{x\in\mathbb{R}^{n}:g(x)<+\infty\}.$
We say that $g$ is proper if $\operatorname{dom}g\neq\emptyset$. Given $\rho\geq 0$, a function $g\colon\mathbb{R}^{n}\to\mathbb{\mathbb{R}}\cup\{+\infty\}$ is called $\rho$-weakly convex if $g+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. For $\mu>0$, we say that $g$ is $\mu$-strongly convex if $g-\frac{\mu}{2}\|\cdot\|^{2}$ is convex.
The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
$\operatorname{e}_{\lambda}g(x):=\inf_{y\in\mathbb{R}^{n}}\left(g(y)+\frac{1}{2\lambda}\|x-y\|^{2}\right)\textrm{ for all }x\in\mathbb{R}^{n}.$
If $g$ is proper, lower semicontinuous, and $\rho$-weakly convex, and if $0<\lambda<\frac{1}{\rho}$ (with the convention $1/0=+\infty$), then the above infimum is attained at a unique point $\operatorname{prox}_{\lambda g}(x)\in\mathbb{R}^{n}$, and
$\operatorname{e}_{\lambda}g(x)$ $=g(\operatorname{prox}_{\lambda g}(x))+\frac{1}{2\lambda}\|x-\operatorname{prox}_{\lambda g}(x)\|^{2}.$
In this case, the operator $x\mapsto\operatorname{prox}_{\lambda g}(x)$ is everywhere defined and is called the proximal operator of $g$ of index $\lambda$. Moreover, $\operatorname{e}_{\lambda}g$ is continuously differentiable and, for each $x\in\mathbb{R}^{n}$,
$\nabla\operatorname{e}_{\lambda}g(x)=\frac{1}{\lambda}(x-\operatorname{prox}_{\lambda g}(x)).$
We refer to [1, 2] for more details.
Let $\delta>0$, and let $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a proper function such that $e^{-g/\delta}$ is integrable and
$\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy>0.$
Under these assumptions, $e^{-g/\delta}$ induces a probability measure. We denote by
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}$
the expectation of a given function $w$ with respect to the probability measure $\sigma_{\delta}$, which is formally defined for every Lebesgue-measurable set $A$ as
$\sigma_{\delta}(A):=\frac{\int_{A}e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
3. Rates of Convergence for Nonsmooth Functions
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The goal of this section is to establish an explicit convergence rate for the map
$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[y\,\exp(-f(y)/\delta)\bigr]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[\exp(-f(y)/\delta)\bigr]},$
and to quantify its approximation of the proximal point:
$m_{\delta}(x)\;\longrightarrow\;\operatorname{prox}_{\lambda f}(x)\qquad\text{as }\delta\downarrow 0.$
It is straightforward to verify that if $f$ is $\rho$-weakly convex for some $\rho\geq 0$ and $\operatorname{dom}f$ has nonempty interior, then $m_{\delta}$ is well-defined for every $\delta>0$ and $0<\lambda<1/\rho$.
The following result provides a convergence rate for this barycenter toward the proximal point.
Theorem 3.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be proper, lower semicontinuous, and $\rho$-weakly convex for some $\rho\geq 0$. Assume that $\operatorname{dom}f$ has nonempty interior. Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\sqrt{\frac{n\delta}{\mu}}.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex, it follows that $g$ is $\mu$-strongly convex and
(2) $g(y)\geq g(\operatorname{prox}_{\lambda f}(x))+\frac{\mu}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\textrm{ for all }y\in\mathbb{R}^{n}.$
Consider the vector field
$F(y):=(y-\operatorname{prox}_{\lambda f}(x))e^{-g(y)/\delta}.$
Now, let us show that $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
Indeed, since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption). On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e. and belongs to $W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D))$. Consequently, $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, and for a.e. $y\in\operatorname{int}(D)$,
$\operatorname{div}F(y)=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}.$
Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$. Since
$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }B_{R}$
and $e^{-g(y)/\delta}\leq e^{-\inf_{B_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
$|\operatorname{div}F(y)|\leq nM_{R,\delta}+\frac{M_{R,\delta}C_{R}}{\delta}\|\nabla g(y)\|,$
Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$. For $R>0$, apply the divergence theorem to $\Omega_{R}$:
$\int_{\Omega_{R}}\operatorname{div}F(y)dy$ $=\int_{\partial\Omega_{R}}F\cdot\nu\,d\mathcal{H}^{n-1}.$
Moreover, by Lemma A.1, we can select $R_{k}\to\infty$ such that
$\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
Now, on the one hand, since $D$ is convex and $\operatorname{prox}_{\lambda f}(x)\in D$, for $\mathcal{H}^{n-1}$-a.e. $y\in\partial D$ one has
$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad F(y)\cdot\nu_{D}(y)\geq 0.$
Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{2}$ is bounded in absolute value by
$CR_{k}^{n}\exp\left(-\frac{\mu}{2\delta}(R_{k}-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }k\to\infty,$
hence the $\partial{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
$0$ $\leq\limsup_{k\to\infty}\int_{B_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\int_{\Omega_{R_{k}}}\operatorname{div}F(y)dy$
$=\int_{D}\operatorname{div}F(y)dy$
$=n\int_{D}e^{-g(y)/\delta}dy-\frac{1}{\delta}\int_{D}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}dy.$
Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the estimate. Next, by strong convexity and almost-everywhere differentiability, we obtain
$\langle\nabla g(y),y-\operatorname{prox}_{\lambda f}(x)\rangle\geq\mu\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}.$
Taking expectation yields
$\mu\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta,$
hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen’s inequality,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}$ $\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|\right]^{2}$
$\leq\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]$
$\leq\frac{n\delta}{\mu}.$
Remark 3.2.
Let $Y\sim\sigma_{\delta}$, where $\sigma_{\delta}$ has density proportional to $\exp(-g/\delta)$. The proof of Theorem 3.1 yields the mean-square localization estimate
$\mathbb{E}\big[\|Y-\operatorname{prox}_{\lambda f}(x)\|^{2}\big]\leq\frac{n\delta}{\mu}.$
This bound also yields a simple concentration inequality: for every $r>0$,
$\sigma_{\delta}\!\big(\|Y-\operatorname{prox}_{\lambda f}(x)\|\geq r\big)\leq\frac{n\delta}{\mu\,r^{2}},$
and therefore, for any $\eta\in(0,1)$,
$\sigma_{\delta}\!\left(B_{\sqrt{n\delta/(\mu\eta)}}(\operatorname{prox}_{\lambda f}(x))\right)\geq 1-\eta.$
The following two examples show that our convergence rates are sharp.
Example 3.1.
Fix $n\geq 1$ and consider
$f(y)=\sum_{i=1}^{n}\max\{y_{i},0\}.$
Then $f$ is proper, lower semicontinuous, convex and $\operatorname{dom}f=\mathbb{R}^{n}$ has nonempty interior. Fix $x=0$ and any $\lambda>0$, and set $\mu=1/\lambda$. Then, $\operatorname{prox}_{\lambda f}(0)=0$ and, by symmetry of $f$, for any $\delta>0$
$m_{\delta}(0)=(a_{\delta},\ldots,a_{\delta}),$
where
$a_{\delta}:=\frac{\int_{\mathbb{R}}t\,e^{-\phi(t)/\delta}\,dt}{\int_{\mathbb{R}}e^{-\phi(t)/\delta}\,dt}\qquad\phi(t):=\max\{t,0\}+\frac{1}{2\lambda}t^{2}.$
Moreover, one can show that
$a_{\delta}=-\sqrt{\frac{2}{\pi}\lambda\delta}+O(\delta)\textrm{ as }\delta\to 0^{+}.$
Consequently,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\sqrt{n}|a_{\delta}|=\sqrt{\frac{2n}{\pi\mu}\delta}+O(\delta).$
This shows that the order of convergence in Theorem 3.1 is sharp.
The following example shows that the bound obtained in Theorem 3.1 is optimal.
Example 3.2.
Fix $n\geq 2$, take $\lambda>0$, set $\mu:=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\},$
where $e_{1}$ denotes the first canonical basis vector. Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]\mathfrak{m}_{\alpha,n},$
where $\chi_{n}$ denotes the $\chi$-distribution with $n$ degrees of freedom and
$\mathfrak{m}_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$. In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $\mathfrak{m}_{\alpha,n}\in[\cos\alpha,1]$ and hence $\mathfrak{m}_{\alpha,n}\to 1$ as $\alpha\searrow 0$. Furthermore,
$\frac{\mathbb{E}[\chi_{n}]}{\sqrt{n}}\to 1\textrm{ as }n\to\infty.$
Consequently, for every $\varepsilon>0$ there exist $n$ and $\alpha$ such that, for all $\delta>0$,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
This shows that the rate of convergence in Theorem 3.1 is sharp.
We end this section by showing that $m_{\delta}(x)$ coincides with $\operatorname{prox}_{\lambda f}(x)$ whenever $f$ is quadratic.
Example 3.3.
Let $A$ be a symmetric and satisfy $A\succeq-\rho I$ for some $\rho\geq 0$, and define
$f(y)=\frac{1}{2}\langle Ay,y\rangle+\langle b,y\rangle+c.$
Then for every $x\in\mathbb{R}^{n}$, every $\delta>0$, and every $0<\lambda<1/\rho$ (with the convention $1/0=+\infty$), one has
$m_{\delta}(x)=\operatorname{prox}_{\lambda f}(x)=\left(A+\frac{1}{\lambda}I\right)^{-1}\left(\frac{1}{\lambda}x-b\right).$
3.1. Stochastic Approximation of Projections onto Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. We define
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
As a direct consequence of Theorem 3.1, we obtain a quantified convergence result: $p_{\delta}(x)\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$.
Corollary 3.3.
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Fix $x\in\mathbb{R}^{n}$. Then, for every $\delta>0$,
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq\sqrt{n\,\delta}.$
4. Smooth results
In this section, we show that additional smoothness of the problem data yields improved convergence rates compared with those in Section 3.
Theorem 4.1.
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}$ be $C^{2}$ and $\rho$-weakly convex with $\rho\geq 0$. Assume that $\nabla^{2}f$ is globally Lipschitz: there exists $L\geq 0$ such that
$\|\nabla^{2}f(u)-\nabla^{2}f(v)\|\leq L\|u-v\|\quad\textrm{ for all }u,v\in\mathbb{R}^{n}.$
Fix $x\in\mathbb{R}^{n}$ and $0<\lambda<1/\rho$, and set $\mu:=\frac{1}{\lambda}-\rho>0$. Then, for every $\delta>0$,
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta.$
Proof.
Fix $x\in\mathbb{R}^{n}$ and assume that $0<\lambda<1/\rho$. Let us consider
$g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}.$
Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and that (2) holds. In particular, (2) implies that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
Given a function $w(\cdot)$, recall that
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
Claim 1: $\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=0$.
Proof of Claim 1: First, note that the Lipschitz continuity of the Hessian of $f$ implies that $\nabla g$ grows at most quadratically. More precisely, there exist constants $a,b>0$ such that
$\|\nabla g(y)\|\leq a\|y\|^{2}+b\qquad\text{for all }y\in\mathbb{R}^{n}.$
In particular, $\nabla g$ is integrable with respect to the probability measure $\sigma_{\delta}$.
Second, we observe that
$\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)=-\frac{1}{\delta}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}.$
Moreover, from (2), it follows that
$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{2\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
which implies that
$0=\int_{\mathbb{R}^{n}}\frac{\partial}{\partial y_{i}}\left(e^{-g(y)/\delta}\right)dy=-\frac{1}{\delta}\int_{\mathbb{R}^{n}}\frac{\partial g}{\partial y_{i}}(y)e^{-g(y)/\delta}dy,$
which implies the claim.
Claim 2: Let $H=\nabla^{2}g(\operatorname{prox}_{\lambda f}(x))$. Then, for all $y\in\mathbb{R}^{n}$
$\nabla g(y)=H(y-\operatorname{prox}_{\lambda f}(x))+r(y),$
where $\|r(y)\|\leq\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}$.
Proof of Claim 2: Since $\nabla g(\operatorname{prox}_{\lambda f}(x))=0$, the formula follows from Taylor’s theorem with
$r(y)=\int_{0}^{1}(\nabla^{2}g(\operatorname{prox}_{\lambda f}(x)+t(y-\operatorname{prox}_{\lambda f}(x)))-H)dt.$
Because $\nabla^{2}g=\nabla^{2}f+\frac{1}{\lambda}I$ and $\nabla^{2}f$ is $L$-Lipschitz, we have
$\|\nabla^{2}g(u)-\nabla^{2}g(v)\|\leq L\|u-v\|.$
Hence
$\|r(y)\|\leq\int_{0}^{1}Lt\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}=\frac{L}{2}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2},$
which proves the claim.
Claim 3: $m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)=-H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]$.
Proof of Claim 3: From Claim 1 and 2 we have that
$0$ $=\mathbb{E}_{\sigma_{\delta}}[\nabla g(y)]=H(m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x))+\mathbb{E}_{\sigma_{\delta}}[r(y)],$
which implies the result.
Claim 4: For every $\delta>0$, we have
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\delta.$
Proof of Claim 4: On the one hand, from Claims 2 and 3, we have
(3) $\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|$ $=\|H^{-1}\mathbb{E}_{\sigma_{\delta}}[r(y)]\|$
$\leq\frac{1}{\mu}\mathbb{E}_{\sigma_{\delta}}\left[\|r(y)\|\right]$
$\leq\frac{L}{2\mu}\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}],$
where we used that $H\succcurlyeq\mu I$, hence $\|H^{-1}\|_{2}\leq 1/\mu$, where $\|\cdot\|_{2}$ is the induced $\ell_{2}$-operator norm. Moreover, using that $\mathbb{E}_{\sigma_{\delta}}[y]=m_{\delta}(x)$, we get that
(4) $\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]=\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]+\mathbb{E}_{\sigma_{\delta}}[\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}].$
On the other hand, we observe that $g/\delta$ is $\mu/\delta$ strongly convex. Hence, $h_{\delta}=e^{-g/\delta}$ is strongly log-concave. From Brascamp–Lieb inequality (see [3, Theorem 4.1]), we get that
(5) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\qquad\Rightarrow\qquad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu},$
where $\operatorname{Cov}_{\sigma_{\delta}}$ denotes the covariance matrix of a random vector under the probability measure $\sigma_{\delta}$. Moreover, Theorem 3.1 gives us the estimation
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{n\delta}{\mu}.$
Plugging (5) and the above inequality into (4) yields
$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{n\delta}{\mu}=\frac{2n\delta}{\mu}.$
Finally, from the above inequality and (3), we get
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{2n\delta}{\mu}=\frac{nL}{\mu^{2}}\delta,$
which proves Claim 4 and the theorem. ∎
4.1. Stochastic Approximation of Projections onto Smooth Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed set with nonempty interior. Recall that
$p_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\cdot\mathbf{1}_{C}(y)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[\mathbf{1}_{C}(y)]}=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}[y\mid y\in C].$
Definition 4.2.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ (see Figure 1) if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
$C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)=\{(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}:\ t\leq h(z)\},$
and $\nabla^{2}h$ is Lipschitz on $\mathbb{B}^{n-1}_{\rho}$ with constant $M\geq 0$, i.e.,
$\|\nabla^{2}h(z)-\nabla^{2}h(z^{\prime})\|\leq M\|z-z^{\prime}\|\qquad\text{for all }z,z^{\prime}\in\mathbb{B}^{n-1}_{\rho}.$
Figure 1. Local $C^{2,1}$ chart of $\partial C$.
In the case of a smooth convex set, we obtain the following refinement of Corollary 3.3, which yields a better convergence rate of order $\delta$.
Theorem 4.3.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exists $\delta_{0}>0$ such that for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=$ $\mathcal{O}(\delta).$
Proof.
The proof can be found in Section B. ∎
Remark 4.4.
A careful review of the proof of Theorem 4.3 shows that, for this result, it suffices that the set $C$ be locally a convex body in a neighborhood of the (necessarily unique) projection point.
The next example shows that the order of the bound obtained in Theorem 4.3 is sharp.
Example 4.1.
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle\leq 0\}$, where $\|\nu\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,\nu\rangle>0$. Then
$\operatorname{proj}_{C}(x)=x-\langle x,\nu\rangle\nu,\quad d_{C}(x)=\langle x,\nu\rangle.$
Moreover, since $p_{\delta}(x)=\mathbb{E}_{y\sim\mathcal{N}(x,\delta I)}\!\left[\,y\mid y\in C\,\right]$, symmetry implies that the conditional mean in tangential directions is unchanged and the drift is purely normal. Hence,
$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,\nu$
for some scalar $\alpha_{\delta}<0$. In fact, one can show that
$\alpha_{\delta}:=d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})},$
where $\phi$ and $\Phi$ denote the standard normal density and distribution function, respectively. Therefore
$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)\nu,$
which implies that $\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=O(\delta)$ as $\delta\to 0^{+}$. Hence, the order of the bound obtained in Theorem 4.3 is sharp.
Acknowledgements
D. Morales was supported by ANID Chile under the CMM BASAL funds for the Center of Excellence FB210005 and ANID BECAS/DOCTORADO NACIONAL 21252113. P. Pérez-Aros was supported by ANID (Chile) through Fondecyt Regular grants No. 1220886, No. 1240120, and No. 1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17. E. Vilches was supported by ANID (Chile) through Fondecyt Regular grants No. 1220886, No. 1240120, and No. 1261728; CMM BASAL funds for the Center of Excellence FB210005; and the projects ECOS230027 and MATH-AmSud 23-MATH-17.
Appendix A Auxiliary lemmas
Lemma A.1.
Let $D\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and for $R>0$ set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$. Then $\Omega_{R}$ is a bounded Lipschitz domain. Moreover, for a.e. $R>0$,
(6) $\mathcal{H}^{n-1}\big(\partial D\cap\partial B_{R}\big)=0.$
Let $F\in W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$. Then, for every $R>0$ such that (6) holds and $\int_{\Omega_{R}}(|F(x)|+\|\nabla F(x)\|)\,dx<\infty$, we have
$\int_{\Omega_{R}}\operatorname{div}F(x)\,dx=\int_{\partial D\cap B_{R}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}F\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
where $\nu_{D}$ is the outer unit normal to $D$, defined $\mathcal{H}^{n-1}$-a.e. on $\partial D$, $\nu_{B_{R}}$ is the outer unit normal to $B_{R}$. Here, $F$ on $\partial\Omega_{R}$ denotes the interior Sobolev trace of $F|_{\Omega_{R}}$ on $\partial\Omega_{R}$, restricted to $\partial D\cap B_{R}$ and to $\operatorname{int}(D)\cap\partial B_{R}$.
Proof.
The first assertion follows from [7, Chapter 1], since every convex domain has a Lipschitz boundary. To prove (6), we observe that, since $D$ is convex, its boundary $\partial D$ is countably (n1)(n-1)-rectifiable. Hence, there exist countably many Lipschitz maps $\Phi_{k}\colon U_{k}\subset\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ such that
$\mathcal{H}^{n-1}\left(\partial D\setminus\bigcup_{k=1}^{\infty}\Phi_{k}(U_{k})\right)=0.$
see, e.g., [5]. Fix $k$. Define the Lipschitz function $g_{k}\colon U_{k}\to\mathbb{R}$ by $g_{k}(u):=\|\Phi_{k}(u)\|$. By the coarea formula for Lipschitz functions (see [4, Chapter 1]), for a.e. $R>0$ the level set $g_{k}^{-1}(R)$ is (n2)(n-2)-rectifiable and has locally finite $\mathcal{H}^{n-2}$-measure. In particular, $\mathcal{L}^{n-1}(g_{k}^{-1}(R))=0$. Then, using the area estimate for Lipschitz maps (a consequence of the area formula),
$\mathcal{H}^{n-1}(\Phi_{k}(g_{k}^{-1}(R)))\leq\operatorname{Lip}(\Phi_{k})^{n-1}\mathcal{L}^{n-1}(g_{k}^{-1}(R))=0.$
Moreover, $\partial D\cap\Phi_{k}(g_{k}^{-1}(R))\subset\partial D\cap\partial B_{R}$ and $\partial D\cap\partial{B}_{R^{\prime}}\cap\Phi_{k}(g_{k}^{-1}(R))=\emptyset$ for all $R^{\prime}\neq R$. Hence,
$(\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})=(\partial D\cap\partial B_{R})\cap\Phi_{k}(g^{-1}_{k}(R)),$
which justifies that
$\mathcal{H}^{n-1}\big((\partial D\cap\partial B_{R})\cap\Phi_{k}(U_{k})\big)=0.$
Taking the countable union over $k$ and using the $\mathcal{H}^{n-1}$-negligibility of the uncovered part of $\partial D$, we obtain (6). To prove the last assertion, since $\Omega_{R}\subset\operatorname{int}(D)$ is open and bounded, the assumptions on $F$ imply that $F\in W^{1,1}(\Omega_{R};\mathbb{R}^{n})$ and $\operatorname{div}F\in L^{1}(\Omega_{R})$. Since $\Omega_{R}$ is a bounded Lipschitz domain, the (interior) trace operator $\operatorname{Tr}\colon W^{1,1}(\Omega_{R})\to L^{1}(\partial\Omega_{R})$ is well-defined, and the Gauss-Green formula for Sobolev vector fields (see, e.g., [8, Chapter 6]) yields
(7) $\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial\Omega_{R}}\operatorname{Tr}(F)\cdot\nu_{\Omega_{R}}\,d\mathcal{H}^{n-1}.$
Next, we observe that
$\partial\Omega_{R}=(\partial D\cap B_{R})\cup(\operatorname{int}(D)\cap\partial B_{R})\cup(\partial D\cap\partial B_{R}).$
Moreover, on $\partial D\cap B_{R}$ the boundary $\partial\Omega_{R}$ coincides locally with $\partial D$. Hence $\nu_{\Omega_{R}}=\nu_{D}$ $\mathcal{H}^{n-1}$-a.e. on $\partial D\cap B_{R}$. Finally, by (6), the set $\partial D\cap\partial B_{R}$ has $\mathcal{H}^{n-1}$-measure zero and does not contribute to the boundary integral in (7). Therefore (7) splits into
$\int_{\Omega_{R}}\operatorname{div}F\,dx=\int_{\partial D\cap B_{R}}\operatorname{Tr}(F)\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\operatorname{int}(D)\cap\partial B_{R}}\operatorname{Tr}(F)\cdot\nu_{B_{R}}\,d\mathcal{H}^{n-1},$
as claimed. ∎
The following lemma can be proved directly using polar coordinates and integration by parts.
Lemma A.2.
Let $n\geq 1$, $R>0$, $d>0$, and $k\in\{0,1\}$. Then, for every $0<\delta\leq\frac{R^{2}}{2n}$, one has
$\int_{\{y\in\mathbb{R}^{n}:\ \|y\|\geq R\}}(\|y\|+d)^{k}\exp\!\Big(-\frac{\|y\|^{2}}{2\delta}\Big)\,dy\leq\frac{4\pi^{n/2}}{\Gamma\!\left(\frac{n}{2}\right)}(R+d)^{k}R^{n-2}\delta\,\exp\!\Big(-\frac{R^{2}}{2\delta}\Big).$
The following lemma can be found in [6].
Lemma A.3.
For all $\delta>0$ and $y>0$,
$\int_{y}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt\geq\frac{\delta y}{y^{2}+\delta}e^{-\frac{y^{2}}{2\delta}}.$
Appendix B Proof of Theorem 4.3
Step 1: Initialization
Set $p:=\operatorname{proj}_{C}(x)$, $d:=d_{C}(x)$, and $\nu:=\frac{x-p}{d}$. After translating and rotating, we may assume that $p=0$, $\nu=e_{n}$, and $x=de_{n}$. For $y\in\mathbb{R}^{n}$, write $y=(z,t)\in\mathbb{R}^{n-1}\times\mathbb{R}$. Let $H=\nabla^{2}h(0)$ (hence $H\preceq 0$ by concavity, and $\|H\|\leq L$). By Taylor’s theorem with Lipschitz Hessian, for all $z\in\mathbb{B}_{\rho}^{n-1}$.
(8) $h(z)=\frac{1}{2}\langle Hz,z\rangle+r(z),\quad|r(z)|\leq\frac{M}{6}\|z\|^{3},$
and also
(9) $|h(z)|\leq\frac{L}{2}\|z\|^{2}\textrm{ for all }z\in\mathbb{B}_{\rho}^{n-1}.$
Moreover, by convexity of $C$ and the fact that $e_{n}$ is an outward normal at 0 imply that the supporting hyperplane is $\{t=0\}$, hence $h(z)\leq 0$ on $\mathbb{B}_{\rho}^{n-1}$.
Define (see Figure 2)
$C_{\operatorname{loc}}:=C\cap\big(\mathbb{B}^{n-1}_{\rho}\times(-\rho,\rho)\big)\textrm{ and }C_{\operatorname{tail}}:=C\setminus C_{\operatorname{loc}}.$
Let us define
$N_{\operatorname{loc}}$ $=\int_{C_{\operatorname{loc}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$ $N_{\operatorname{tail}}$ $=\int_{C_{\operatorname{tail}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$
$D_{\operatorname{loc}}$ $=\int_{C_{\operatorname{loc}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy,$ $D_{\operatorname{tail}}$ $=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy.$
With the above definitions, we observe that
$p_{\delta}(x):=\frac{N_{\operatorname{loc}}+N_{\operatorname{tail}}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}.$
Step 2: Bounds on $D_{\operatorname{tail}}$ and $D_{\operatorname{loc}}$.
We observe that $C_{\operatorname{loc}}\subset\mathbb{B}_{\rho}^{n-1}\times(-\rho,0]$ and if $y\in C_{\operatorname{tail}}$, then either $\|z\|\geq\rho$ or $t\leq-\rho$. In both cases,
(10) $\|y-x\|^{2}=\|z\|^{2}+(t-d)^{2}\geq\rho^{2}+d^{2}\textrm{ for }y\in C_{\operatorname{tail}}.$
Figure 2. Local chart of $\partial C$ and the sets $C_{\operatorname{loc}}$ (pale turquoise) and $C_{\operatorname{tail}}$ (light cyan).
Claim 1: Assume that $0<\delta\leq\frac{\rho^{2}+d^{2}}{2n}$ (with the convention that $1/0=+\infty$). Then
$D_{\operatorname{tail}}:=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\leq\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)}(\rho^{2}+d^{2})^{\frac{n-2}{2}}\delta e^{-\frac{\rho^{2}+d^{2}}{2\delta}}.$
Proof of Claim 1: Since $C_{\operatorname{tail}}\subset\{y\in\mathbb{R}^{n}:\|y-x\|^{2}\geq\rho^{2}+d^{2}\}$, we have
$D_{\operatorname{tail}}=\int_{C_{\operatorname{tail}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\leq\int_{\{y\in\mathbb{R}^{n}:\|y-x\|^{2}\geq\rho^{2}+d^{2}\}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy.$
Applying Lemma A.2 with $R=\sqrt{\rho^{2}+d^{2}}$ yields the claim. ∎
Claim 2: Assume that $0<\delta\leq\min\left\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}}\right\}$.
Then
$D_{\operatorname{loc}}:=\int_{C_{\operatorname{loc}}}e^{-\frac{\|y-x\|^{2}}{2\delta}}dy\geq\left[\psi(\delta)\cdot\mathcal{K}_{\operatorname{err}}(\delta)\cdot\frac{C_{\operatorname{vol}}(n)}{(1+dL)^{\frac{n-1}{2}}}\right]\cdot\delta^{\frac{n-1}{2}}\cdot e^{-\frac{d^{2}}{2\delta}},$
where
$\psi(\delta)$ $=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}\omega(z),\quad r_{\delta}=\sqrt{\frac{\delta}{1+dL}},$
$\omega(z)$ $=\frac{\delta(d-h(z))}{(d-h(z))^{2}+\delta}-\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-(d-h(z))^{2}\right]}{2\delta}\right),$
$\mathcal{K}_{\operatorname{err}}(\delta)$ $=\exp\left(-\frac{L^{2}\delta}{8(1+dL)^{2}}\right),$
$C_{\operatorname{vol}}(n)$ $=\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv.$
Proof of Claim 2: Since $\delta\leq\rho^{2}(1+dL)$, we have that $r_{\delta}\leq\rho$. Set A(z):=dh(z)A(z):=d-h(z) and observe that
  • By (9), we have
    $A(z)=d-h(z)\leq d+\frac{L}{2}r_{\delta}^{2}=d+\frac{L\delta}{2(1+dL)}=:A_{\max}.$
  • Since $0<\delta<\frac{2\rho(1+dL)}{L}$, it follows that $A_{\max}<d+\rho$. Consequently,
    $(d+\rho)^{2}-A(z)^{2}\geq(d+\rho)^{2}-A_{\max}^{2}.$
  • Since $0<\delta<\frac{d\rho}{1+\frac{dL}{2(1+dL)}}$, we obtain the lower bound
    $J(z)$ $=\int_{A(z)}^{d+\rho}e^{-\frac{t^{2}}{2\delta}}\,dt$
    $=\int_{A(z)}^{+\infty}e^{-\frac{t^{2}}{2\delta}}\,dt-\int_{\rho+d}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
    $\geq\frac{\delta A(z)}{A(z)^{2}+\delta}e^{-\frac{A(z)^{2}}{2\delta}}-\frac{\delta}{\rho+d}e^{-\frac{(\rho+d)^{2}}{2\delta}}$
    $=e^{-\frac{A(z)^{2}}{2\delta}}\omega(z)$
    $\geq\psi(\delta)\,e^{-\frac{A(z)^{2}}{2\delta}},$
    where we have used Lemma A.3.
Therefore,
$D_{\operatorname{loc}}$ $=\int_{z\in\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\left(\int_{-\rho}^{h(z)}e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)\,dz$
$=\int_{z\in\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}J(z)\,dz$
$\geq\int_{z\in\mathbb{B}_{r_{\delta}}^{n-1}}\psi(\delta)e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz.$
Define $E(z):=\|z\|^{2}+A(z)^{2}$. Since $\|z\|\leq r_{\delta}$, it follows that
$E(z)\leq d^{2}+(1+dL)\|z\|^{2}+\frac{L^{2}}{4}\|z\|^{4}\leq d^{2}+(1+dL)\|z\|^{2}+\frac{L^{2}}{4}\left(\frac{\delta}{1+dL}\right)^{2}.$
Hence,
$D_{\operatorname{loc}}$ $\geq\psi(\delta)e^{-\frac{L^{2}\delta}{8(1+dL)^{2}}}e^{-\frac{d^{2}}{2\delta}}\left(\frac{\delta}{1+dL}\right)^{\frac{n-1}{2}}\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv$
$=\left[\psi(\delta)\cdot\mathcal{K}_{\operatorname{err}}(\delta)\cdot\frac{C_{\operatorname{vol}}(n)}{(1+dL)^{\frac{n-1}{2}}}\right]\cdot\delta^{\frac{n-1}{2}}\cdot e^{-\frac{d^{2}}{2\delta}},$
which proves the claim.
Claim 3: If $0<\delta\leq\delta_{0}:=\min\big\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}},\frac{\rho^{2}+d^{2}}{2n}\big\}$, then $\psi(\delta)>0$ and
$\frac{D_{\operatorname{tail}}}{D_{\operatorname{loc}}}\leq\left[\frac{\mathcal{K}_{\operatorname{geom}}}{\mathcal{K}_{\operatorname{err}}(\delta)}\right]\left(\frac{\delta}{\psi(\delta)}\right)\delta^{-\frac{(n-1)}{2}}e^{-\frac{\rho^{2}}{2\delta}},$
where $\mathcal{K}_{\operatorname{err}}(\delta)$ and $\psi(\delta)$ are given in Claim 2. Moreover,
$\mathcal{K}_{\operatorname{geom}}$ $=\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)C_{\operatorname{vol}}(n)}(\rho^{2}+d^{2})^{\frac{n-2}{2}}(1+\rho L)^{\frac{n-1}{2}}.$
Proof of Claim 3: We observe that for $0<\delta\leq\delta_{0}$,
$\omega(z)$ $=\frac{\delta A(z)}{A(z)^{2}+\delta}-\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-A(z)^{2}\right]}{2\delta}\right)$
$\geq\delta\left(\frac{A_{\operatorname{max}}}{A_{\operatorname{max}}^{2}+\delta}-\frac{1}{e(\rho+d)}\right)>0,$
which implies that $\psi(\delta)=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}\omega(z)>0$. Finally, the result follows from Claims 1 and 2. ∎
Claim 4: Let $\delta_{1}:=\min\left\{\delta_{0},\frac{d^{2}}{3},\frac{\rho^{2}}{2\ln\!\left(4\left(1+\frac{\rho}{d}\right)\right)}\right\}$. Then, for all $\delta\in(0,\delta_{1})$, one has
$\psi(\delta)\geq\frac{\delta}{2d}.$
Proof of Claim 4: From the definition of $\omega(z)=T_{1}(z)-T_{2}(z)$, where
$T_{1}(z)=\frac{\delta A(z)}{A(z)^{2}+\delta}\quad\textrm{ and }\quad T_{2}(z)=\frac{\delta}{\rho+d}\exp\left(-\frac{\left[(\rho+d)^{2}-A(z)^{2}\right]}{2\delta}\right)$
Hence, on the one hand, if $\delta\leq d^{2}/3$, then $T_{1}(z)\geq\frac{3\delta}{4d}$. On the other hand, if $\delta\leq\frac{\rho^{2}}{2\ln\left(4(1+\frac{\rho}{d})\right)}$, then $T_{2}(z)\leq\frac{\delta}{4d}$. Finally,
$\psi(\delta)=\inf_{z\in\mathbb{B}^{n-1}_{r_{\delta}}}(T_{1}(z)-T_{2}(z))\geq\frac{3\delta}{4d}-\frac{\delta}{4d}=\frac{\delta}{2d}.$
With respect to the tangential/normal splitting induced by $y=(z,t)$, we write $N_{\operatorname{loc}}=(N_{\operatorname{loc}}^{\tau},N_{\operatorname{loc}}^{\nu})$ and $N_{\operatorname{tail}}=(N_{\operatorname{tail}}^{\tau},N_{\operatorname{tail}}^{\nu})$. Hence,
$N_{\operatorname{loc}}^{\tau}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}\left(\int_{-\rho}^{h(z)}e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)\,dz$
$N_{\operatorname{loc}}^{\nu}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\underbrace{\left(\int_{-\rho}^{h(z)}t\,e^{-\frac{(t-d)^{2}}{2\delta}}\,dt\right)}_{=:I_{\nu}(z)}\,dz$
Moreover, we write $p_{\delta}(x)=(p_{\delta}^{\tau}(x),p_{\delta}^{\nu}(x))$, where
(11) $p_{\delta}^{\tau}(x)=\frac{N_{\operatorname{loc}}^{\tau}+N_{\operatorname{tail}}^{\tau}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\quad p_{\delta}^{\nu}(x)=\frac{N_{\operatorname{loc}}^{\nu}+N_{\operatorname{tail}}^{\nu}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}.$
Step 3: Bounding $N_{\operatorname{tail}}$.
Claim 5:
For all $0<\delta\leq\frac{d^{2}+\rho^{2}}{2n}$, one has
$\|N_{\operatorname{tail}}\|\leq\frac{4\pi^{n/2}}{\Gamma\left(\frac{n}{2}\right)}\sqrt{\rho^{2}+d^{2}}(\rho^{2}+d^{2})^{\frac{n-2}{2}}\delta e^{-\frac{\rho^{2}+d^{2}}{2\delta}}.$
Proof of Claim 5: From (10), one has
$\|N_{\operatorname{tail}}\|$ $=\left\|\int_{C_{\operatorname{tail}}}ye^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy\right\|$
$\leq\int_{C_{\operatorname{tail}}}\|y\|e^{-\frac{\|y-x\|^{2}}{2\delta}}\,dy$
$\leq\int_{\{u\in\mathbb{R}^{n}:\|u\|\geq\sqrt{\rho^{2}+d^{2}}\}}(\|u\|+d)e^{-\frac{\|u\|^{2}}{2\delta}}du,$
which, by virtue of Lemma A.2, implies the claim.
Step 4: Bounding the normal component $N_{\operatorname{loc}}^{\nu}$ of $N_{\operatorname{loc}}$.
Claim 6:
The following formula holds:
(12) $N_{\operatorname{loc}}^{\nu}=d\,D_{\operatorname{loc}}-\delta I_{\operatorname{num}}+\delta\,e^{-\frac{(\rho+d)^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz,$
where $I_{\operatorname{num}}:=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz$.
Proof of Claim 6: Recall that $N_{\operatorname{loc}}^{\nu}=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}I_{\nu}(z)\,dz$ with
$I_{\nu}(z)=\int_{-\rho}^{h(z)}t\,e^{-\frac{(t-d)^{2}}{2\delta}}\,dt=\underbrace{\int_{-(\rho+d)}^{-A(z)}ue^{-\frac{u^{2}}{2\delta}}\,du}_{\operatorname{Term}_{1}(z)}+d\underbrace{\int_{-(\rho+d)}^{-A(z)}e^{-\frac{u^{2}}{2\delta}}\,du}_{\operatorname{Term}_{2}(z)}.$
Moreover, $\operatorname{Term}_{1}(z)=-\delta\big(e^{-\frac{A(z)^{2}}{2\delta}}-e^{-\frac{(\rho+d)^{2}}{2\delta}}\big)$ and $\operatorname{Term}_{2}(z)=J(z)$. Then,
$I_{\nu}(z)=-\delta e^{-\frac{A(z)^{2}}{2\delta}}+\delta e^{-\frac{(\rho+d)^{2}}{2\delta}}+dJ(z),$
which implies that
$N_{\operatorname{loc}}^{\nu}$ $=\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\left[d\,J(z)-\delta e^{-\frac{A(z)^{2}}{2\delta}}+\delta e^{-\frac{(\rho+d)^{2}}{2\delta}}\right]\,dz$
$=d\,D_{\operatorname{loc}}-\delta I_{\operatorname{num}}+\delta\,e^{-\frac{(\rho+d)^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz,$
which proves the claim. ∎
We observe that the third term in the right-hand side of equality (12) is $\mathcal{O}(e^{-c/\delta})$, $D_{\operatorname{tail}}=\mathcal{O}(\delta e^{-c/\delta})$ (see Claim 1), and $\|N_{\operatorname{tail}}\|=\mathcal{O}(\delta e^{-c/\delta})$ (see Claim 5). Hence, informally speaking, by Claim 6, we get
$p_{\delta}^{\nu}(x)=\frac{N_{\operatorname{loc}}^{\nu}+N_{\operatorname{tail}}^{\nu}}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\approx\frac{N_{\operatorname{loc}}^{\nu}}{D_{\operatorname{loc}}}\approx d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}.$
Therefore, we consider the estimator
$\tilde{p}_{\delta}^{\nu}:=d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}.$
Let us consider the measure $\eta_{\delta}$ defined on $\mathbb{B}^{n-1}_{\rho}$ whose density is proportional to
$e^{-\frac{\|z\|^{2}+(d-h(z))^{2}}{2\delta}}.$
Claim 7: Assume that $0<\delta\leq\min\left\{\rho^{2}(1+dL),\frac{2\rho(1+dL)}{L},\frac{d\rho}{1+\frac{dL}{2(1+dL)}}\right\}$. Then there exists a function $r(A(z),\delta)$ such that $0<r(A(z),\delta)<\frac{\delta}{A(z)^{2}}$ and
$\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}=\frac{1}{\delta\mathbb{E}_{\eta_{\delta}}\left[\frac{1-r(A(z),\delta)}{A(z)}\right]-\Delta(\delta)}\quad\textrm{ for all }0<\delta\leq\delta_{1},$
where
$0\leq\Delta(\delta)\leq\frac{(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\operatorname{err}}\cdot(\rho+d)}\rho^{n-1}\frac{\operatorname{Vol}(\mathbb{B}^{n-1})}{\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv}\cdot\delta^{\frac{3-n}{2}}e^{-\frac{\rho^{2}+2d\rho}{2\delta}}=\mathcal{O}(\delta^{\frac{3-n}{2}}e^{-c/\delta}).$
Proof of Claim 7: Let us observe that
$J(z)$ $=\int_{A(z)}^{d+\rho}e^{-\frac{t^{2}}{2\delta}}\,dt$
$=\int_{A(z)}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt-\int_{d+\rho}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
$=\frac{\delta}{A(z)}e^{-\frac{A(z)^{2}}{2\delta}}\big(1-\underbrace{A(z)e^{\frac{A(z)^{2}}{2\delta}}\int_{A(z)}^{\infty}\frac{1}{t^{2}}e^{-\frac{t^{2}}{2\delta}}\,dt}_{=:r(A(z),\delta)}\big)-\int_{d+\rho}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt.$
Moreover, since $t\geq A(z)$ on $A(z),\infty)$, we have $1/t^{2}\leq 1/A(z)^{2}$. Using this bound and the standard Gaussian tail estimate (obtained by integration by parts), for all $w>0$
(13) $\int_{w}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt\leq\frac{\delta}{w}e^{-\frac{w^{2}}{2\delta}},$
we obtain
$0<r(A(z),\delta)$ $=A(z)e^{\frac{A(z)^{2}}{2\delta}}\int_{A(z)}^{\infty}\frac{1}{t^{2}}e^{-\frac{t^{2}}{2\delta}}\,dt$
$\leq A(z)e^{\frac{A(z)^{2}}{2\delta}}\cdot\frac{1}{A(z)^{2}}\int_{A(z)}^{\infty}e^{-\frac{t^{2}}{2\delta}}\,dt$
$\leq A(z)e^{\frac{A(z)^{2}}{2\delta}}\cdot\frac{1}{A(z)^{2}}\left(\frac{\delta}{A(z)}e^{-\frac{A(z)^{2}}{2\delta}}\right)$
$=\frac{\delta}{A(z)^{2}}.$
Therefore,
$\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}$ $=\frac{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz}{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}}{2\delta}}J(z)\,dz}$
$=\frac{1}{\delta\mathbb{E}_{\eta_{\delta}}\left[\frac{1-r(A(z),\delta)}{A(z)}\right]-\Delta(\delta)},$
where
$\Delta(\delta):=\frac{\int_{\rho+d}^{\infty}e^{-\frac{u^{2}}{2\delta}}\,du\cdot\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}}{2\delta}}\,dz}{\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz}.$
On the one hand, by inequality (13), we observe that the numerator of $\Delta(\delta)$ is bounded by
$\frac{\delta}{\rho+d}e^{-\frac{(\rho+d)^{2}}{2\delta}}\operatorname{Vol}(\mathbb{B}^{n-1}_{\rho}).$
On the other hand, the analysis for the denominator is similar to the analysis of $D_{\operatorname{loc}}$ (see Claim 2). Hence,
$\int_{\mathbb{B}^{n-1}_{\rho}}e^{-\frac{\|z\|^{2}+A(z)^{2}}{2\delta}}\,dz\geq\mathcal{K}_{\operatorname{err}}\cdot e^{-\frac{d^{2}}{2\delta}}\left(\frac{\delta}{1+dL}\right)^{\frac{n-1}{2}}\int_{\mathbb{B}^{n-1}}e^{-\frac{\|v\|^{2}}{2}}dv.$
Combining the above inequalities, we get the claim. ∎
Claim 8: We have
$\tilde{p}_{\delta}^{\nu}=\mathbb{E}_{\eta_{\delta}}[h(z)]-d\,\mathbb{E}_{\eta_{\delta}}\!\left[r(A(z),\delta)\right]+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
Moreover,
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2})\textrm{ as }\delta\to 0^{+}.$
Proof of Claim 8: Let $g_{\delta}(z)=\frac{1-r(A(z),\delta)}{A(z)}$. Then, by virtue of Claim 7, we get
$\tilde{p}_{\delta}^{\nu}$ $=d-\delta\frac{I_{\operatorname{num}}}{D_{\operatorname{loc}}}=d-\frac{\delta}{\delta\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}+O(\delta^{\frac{1-n}{2}}e^{-c/\delta}).$
Moreover, by Taylor’s theorem,
$\frac{1}{\mathbb{E}_{\eta_{\delta}}[g_{\delta}(z)]}=\frac{1}{\frac{1}{d}+\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]}=d-d^{2}\mathbb{E}_{\eta_{\delta}}\left[g_{\delta}(z)-\frac{1}{d}\right]+O(\delta^{2}).$
Therefore,
$\tilde{p}_{\delta}^{\nu}=\mathbb{E}_{\eta_{\delta}}[h(z)]-d\,\mathbb{E}_{\eta_{\delta}}[r(A(z),\delta)]+O(\delta^{2}).$
Finally, using that $|h(z)|\leq\frac{L}{2}\|z\|^{2}$ and $\mathbb{E}_{\eta_{\delta}}[\|z\|^{2}]\leq(n-1)\delta+O(\delta^{3/2})$, we obtain that
$|\tilde{p}_{\delta}^{\nu}|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}),$
which proves the claim.
Step 4: Bounding the tangential component $N_{\operatorname{loc}}^{\tau}$ of $N_{\operatorname{loc}}$.
Claim 9:
For all $0<\delta\leq\delta_{1}$
$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{M(n^{2}-1)}{6}(2\pi\delta)^{\frac{n-1}{2}}\delta^{2}e^{-\frac{d^{2}}{2\delta}}.$
Proof of Claim 9: Recall that
$N_{\operatorname{loc}}^{\tau}=\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}J(z)dz.$
Since $\mathbb{B}_{\rho}^{n-1}$ is symmetric and $e^{-\frac{\|z\|^{2}}{2\delta}}$ is even,
$N_{\operatorname{loc}}^{\tau}=\frac{1}{2}\int_{\mathbb{B}_{\rho}^{n-1}}ze^{-\frac{\|z\|^{2}}{2\delta}}(J(z)-J(-z))dz.$
Hence
$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{1}{2}\int_{\mathbb{B}_{\rho}^{n-1}}\|z\|e^{-\frac{\|z\|^{2}}{2\delta}}|J(z)-J(-z)|dz.$
Set $F(u):=\int_{-\rho}^{u}e^{-\frac{(t-d)^{2}}{2\delta}}dt$. Hence, $J(z)=F(h(z))$ and $F^{\prime}(u)=e^{-\frac{(u-d)^{2}}{2\delta}}$. By the mean value theorem, for each $z$ there exists $\theta_{z}$ between $h(z)$ and h(z)h(-z) such that
$J(z)-J(-z)=F^{\prime}(\theta_{z})(h(z)-h(-z)).$
Moreover, since $h(z)\leq 0$ and $h(-z)\leq 0$, we have $\theta_{z}\leq 0$, hence $(\theta_{z}-d)^{2}\geq d^{2}$ and thus
$|F^{\prime}(\theta_{z})|\leq e^{-\frac{d^{2}}{2\delta}}.$
Therefore,
$|J(z)-J(-z)|\leq e^{-\frac{d^{2}}{2\delta}}|h(z)-h(-z)|.$
From (8), we get that
$h(z)=\frac{1}{2}\langle Hz,z\rangle+r(z),\quad|r(z)|\leq\frac{M}{6}\|z\|^{3}.$
Since the quadratic term is even,
h(z)h(z)=$r(z)$,h(z)-h(-z)=r(z)-r(-z),
and hence $|J(z)-J(-z)|\leq\frac{M}{3}\|z\|^{3}e^{-\frac{d^{2}}{2\delta}}$. Using the above estimates, we get
$\|N_{\operatorname{loc}}^{\tau}\|\leq\frac{M}{6}e^{-\frac{d^{2}}{2\delta}}\int_{\mathbb{B}_{\rho}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz\leq\frac{M}{6}e^{-\frac{d^{2}}{2\delta}}\int_{\mathbb{R}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz$
If $W\sim\mathcal{N}(0,\delta I_{n-1})$, then
$\int_{\mathbb{R}^{n-1}}\|z\|^{4}e^{-\frac{\|z\|^{2}}{2\delta}}dz=(2\pi\delta)^{\frac{n-1}{2}}\mathbb{E}\|W\|^{4},\quad\mathbb{E}\|W\|^{4}=\delta^{2}(n^{2}-1),$
which yields the claim. ∎
Claim 10: Under the assumptions of Theorem 4.3, we have
$\|p_{\delta}^{\tau}(x)\|\leq\frac{M(n^{2}-1)d}{6}\delta+O(\delta^{2})\textrm{ as }\delta\to 0^{+}.$
Proof of Claim 10: By (11),
$\|p_{\delta}^{\tau}(x)\|\leq\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}+\frac{\|N_{\operatorname{tail}}\|}{D_{\operatorname{loc}}+D_{\operatorname{tail}}}\leq\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}}+\frac{\|N_{\operatorname{tail}}\|}{D_{\operatorname{loc}}}.$
Using Claim 9 and Claim 2-4, for $0<\delta\leq\delta_{1}$ we have
$\frac{\|N_{\operatorname{loc}}^{\tau}\|}{D_{\operatorname{loc}}}$ $\leq\frac{\frac{M(n^{2}-1)}{6}(2\pi\delta)^{\frac{n-1}{2}}\delta^{2}e^{-\frac{d^{2}}{2\delta}}}{\psi(\delta)\mathcal{K}_{\rm err}(\delta)\frac{C_{\rm vol}(n)}{(1+dL)^{\frac{n-1}{2}}}\delta^{\frac{n-1}{2}}e^{-\frac{d^{2}}{2\delta}}}$
$=\frac{M(n^{2}-1)}{6}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\rm err}(\delta)C_{\rm vol}(n)}\cdot\frac{\delta^{2}}{\psi(\delta)}.$
By Claim 4, $\psi(\delta)\geq\delta/(2d)$, hence
$\frac{\|N_{\rm loc}^{\tau}\|}{D_{\rm loc}}\leq\frac{M(n^{2}-1)d}{3}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{\mathcal{K}_{\rm err}(\delta)C_{\rm vol}(n)}\cdot\delta.$
Finally, Claim 5 and Claim 2–4 imply $\|N_{\rm tail}\|/D_{\rm loc}=O(e^{-c/\delta})$, which can be absorbed into $O(\delta^{2})$ as $\delta\to 0^{+}$. This completes the proof. ∎
Step 5: Conclusion.
Since $\operatorname{proj}_{C}(x)=0$ in the chosen coordinates, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|=\|p_{\delta}(x)\|\leq|p_{\delta}^{\nu}(x)|+\|p_{\delta}^{\tau}(x)\|.$
By Claim 8 (together with Claims 1,5,6 to control the tail and the exponentially small terms), we obtain
$|p_{\delta}^{\nu}(x)|\leq\left(\frac{L(n-1)}{2}+\frac{1}{d}\right)\delta+O(\delta^{3/2}).$
By Claim 10,
$\|p_{\delta}^{\tau}(x)\|\leq C_{\tau}\,\delta+O(\delta^{2}),\qquad C_{\tau}=\frac{M(n^{2}-1)d}{3}\,\frac{(2\pi)^{\frac{n-1}{2}}(1+dL)^{\frac{n-1}{2}}}{C_{\rm vol}(n)}.$
Combining the two estimates yields the theorem. ∎